OpenStudy (anonymous):

Solve for v in the formula v = u + 10t when u = 16 and t = 4.

OpenStudy (anonymous):

Substituting the values of u and t in the given equation, we get v = 16 + 10(4) = 16 + 40 = 56

OpenStudy (alekos):

Are you serious?